### v1

# Enhancing Facial Expression Recognition through Dual-Direction Attention Mixed Feature Networks: Application to 7th ABAW Challenge

Abstract
We present our contribution to the 7th ABAW challenge at ECCV 2024, by utilizing a Dual-Direction Attention Mixed Feature Network for multitask facial expression recognition we achieve results far beyond the proposed baseline for the Multi-Task ABAW challenge. Our proposal uses the well-known DDAMFN architecture as base to effectively predict valence-arousal, emotion recognition, and action units. We demonstrate the architecture ability to handle these tasks simultaneously, providing insights into its architecture and the rationale behind its design. Additionally, we compare our results for a multitask solution with independent single-task performance.
Keywords:
ABAW Challenge Emotion recogntion DDAMFN
1 Introduction
Facial emotion recognition has emerged as a pivotal area of research within affective computing, driven by its potential applications in fields ranging from human-computer interaction to psychological research and clinical diagnostics. Since Ekman’s classification of human expression faces into emotions [4], a lot of studies have emerged, in recent years: Calvo et al. [2], Baltrusaities et al. [1], or Kaya et al. [5] laid foundational frameworks for understanding facial expressions as a window into emotional states. In contemporary research, [20, 6] continued to refine and expand these methodologies. By synthesizing insights from cognitive psychology, computer vision, and machine learning, researchers have made significant strides in enhancing the accuracy and applicability of facial emotion recognition systems. Moreover, the integration of valence and arousal dimensions [25, 26] adds depth to the interpretation of emotional states, enabling more nuanced insights into human affective experiences.
Action unit detection [3, 21] complements these efforts by parsing facial expressions into discrete muscle movements, facilitating a finer-grained analysis of emotional expressions across cultures and contexts. Such advancements not only improve the reliability of automated emotion recognition systems but also open avenues for personalized affective computing applications in fields such as mental health monitoring [22] and user experience design [27].
To tackle these challenges, researchers have explored innovative architectures such as the Dual-Direction Attention Mixed Feature Network (DDAMFN) [28]. This approach integrates attention mechanisms and mixed feature extraction, enhancing the network’s ability to capture intricate details within facial expressions. This architecture shows promising results in multitask challenges, together with other pretrained networks [24].
There exists a modern day need to create machines capable to comprehend and appropriately respond to human feelings day-to-day on-the-wild applications. This challenge was presented as a form of one of the two competitions at the 7th ABAW challenge [13, 10, 9, 8, 18, 17, 12, 7, 16, 15, 11] at ECCV 2024, with a smaller dataset (s-AffWild2) than in previous challenges (AffWild2) [14], but focusing on a multitask scheme [19]. This challenge presents a complex multitask problem that searches for simultaneous predictions of valence-arousal, emotion recognition, and action units. This poses a significant challenge due to the diverse nature of the required outputs. In this work, we respond to the challenge using a DDAMFN-based architecture [28] as a feature extractor. We proposed a fine-tuning training on the s-AffWild2 dataset to increase the performance of the model for a multitask challenge; plus, we evaluated its performance individually on the different task of the challenge: valence-arousal prediction, emotion recognition, and action unit detection.
2 Methodology
2.1 Dataset curation
The s-Aff-Wild2 was introduced by the organizers of 7th ABAW challenge [19] as a subset of Aff-Wild2, representing a subset of image frames of the original videos, without any audio data. Only the frames and the annotations for approximately 221,000 images were presented. The dataset presented a preestablished train-test partition: 142382 instances for training and 26876, for validation. The frames came with annotations in terms of valence-arousal, eight basic expressions (happiness, sadness, anger, fear, surprise, and disgust), the neutral state, and an ’other’ category that included affective states not covered by the aforementioned categories. Additionally, the dataset contained annotations for 12 action units: AU1, AU2, AU4, AU6, AU7, AU10, AU12, AU15, AU23, AU24, AU25, and AU26.
Following the instructions provided by the competition organizers, certain annotation values were clarified to be disregarded. Specifically, annotation values of -5 for valence/arousal, -1 for expressions, and -1 for action units (AUs) were excluded from consideration in the analysis. In response to this directive, the dataset underwent rigorous filtering. Frames containing any annotation values outside the specified acceptable range were excluded from use, even if all other values for that frame were within the normal range. This thorough filtering process ensured that only frames with entirely valid annotations were included, resulting in a refined dataset of 52,154 frames for training and 15,440 frames for validation.
2.2 Network architecture
For the 7th ABAW challenge, we adapted the Dual-Direction Attention Mixed Feature Network (DDAMFN) by replacing the original classifier with a custom one. The custom classifier consists of a dense layer with 512 units followed by three branches, each tailored to a specific task:
  • Valence-Arousal Prediction: 2 output units.
  • Emotion Recognition: 8 output units.
  • Action Units: 12 output units.
Initially, the entire DDAMFN architecture, with weights obtained from training on AffectNet-8, was frozen except for the custom classifier. The classifier was trained in isolation to learn the distinct characteristics of each task without interference from the feature extraction layers. Subsequently, the entire model was unfrozen, and fine-tuning was performed on all layers. This comprehensive training stage ensured that the feature extraction, attention mechanisms, and classifier were all optimized to work cohesively for the multitask problem.
2.3 Training
The loss for valence-arousal prediction was calculated using the Concordance Correlation Coefficient (CCC). CCC is a measure that evaluates the agreement between two time series by assessing both their precision (how well the observations agree) and accuracy (how well the observations match the true values). For emotion recognition, the loss function used was cross-entropy, which is commonly employed in classification tasks to measure the difference between the predicted probability distribution and the true distribution. For action unit (AU) detection, the binary cross-entropy loss was used, which is suitable for binary classification tasks, measuring the difference between the predicted probability and the actual binary outcome for each action unit. When the whole model is fine-tuned, an extra loss for the attention head is added. This loss works by calculating the mean squared error (MSE) between each pair of attention heads, and then normalizing the total loss by the number of pairs plus a small epsilon to avoid division by zero. This encourages the attention heads to produce consistent outputs.
As a second experiment, we repeated the same procedure. However, this time we applied the method to each task individually. This approach allowed us to compare the results for each specific task, providing more detailed insights and enabling a clearer understanding of the performance variations across different tasks.
Furthermore, for the Action Unit task, a global threshold of 0.5 was initially tested across all AUs, followed by individual optimization for each AU [23].
3 Results
The performance metrics for the multitask challenge are summarized in Table 1. The metrics evaluated include the Concordance Correlation Coefficient (CCC) for valence, arousal, and their combination (valence-arousal), F1 score for emotion classification, F1 score for action unit (AU) detection, and the combined performance score (P Score).
Table 1: Performance metrics for the multitask challenge.
Trainable layers CCC-V CCC-A CCC-V-A F1-Emotion F1-AU P Score
DDAMFN 0.549 0.524 0.536 0.277 0.470 1.287
DDAMFN+thresholds 0.549 0.524 0.536 0.277 0.510 1.327
Custom classifier 0.548 0.518 0.533 0.262 0.473 1.283
Custom classifier+thresholds 0.548 0.518 0.533 0.262 0.500 1.313
Fine-tuning the entire DDAMFN model achieved a CCC of 0.549 for valence, 0.524 for arousal, and 0.536 for valence-arousal. The F1 scores for emotion classification and AU detection were 0.277 and 0.470, respectively, resulting in a P Score of 1.287. Incorporating threshold adjustments slightly improved the F1 score for AU detection to 0.510 and the P Score to 1.327.
Training a custom classifier showed a CCC of 0.548 for valence, 0.518 for arousal, and 0.533 for valence-arousal. The F1 scores for emotion classification and AU detection were 0.262 and 0.473, respectively, with a P Score of 1.283. Threshold adjustments for the custom classifier improved the F1 score for AU detection to 0.500 and the P Score to 1.313.
The performance metrics for individual tasks are detailed in Table 2.
Table 2: Performance metrics for individual tasks.
Trainable layers CCC-V CCC-A CCC-V-A F1-Emotion F1-AU P Score
DDAMFN 0.604 0.550 0.577 0.287 0.490 1.354
DDAMFN+thresholds 0.604 0.550 0.577 0.287 0.529 1.393
Custom classifier 0.530 0.537 0.533 0.243 0.487 1.263
Custom classifier+thresholds 0.530 0.537 0.533 0.243 0.524 1.300
For individual tasks, fine-tuning the entire DDAMFN model demonstrated higher performance with a CCC of 0.604 for valence, 0.550 for arousal, and 0.577 for valence-arousal. The F1 scores for emotion classification and AU detection were 0.287 and 0.490, respectively, with a P Score of 1.354. Threshold adjustments further enhanced the F1 score for AU detection to 0.529 and the P Score to 1.393.
Training a custom classifier presented a CCC of 0.530 for valence, 0.537 for arousal, and 0.533 for valence-arousal. The F1 scores for emotion classification and AU detection were 0.243 and 0.487, respectively, resulting in a P Score of 1.263. Applying threshold adjustments increased the F1 score for AU detection to 0.524 and the P Score to 1.300.
Overall, both approaches—fine-tuning the entire DDAMFN model and training a custom classifier—performed well across the evaluated metrics, with fine-tuning the DDAMFN model showing slightly better performance in individual task metrics.
4 Conclusion
The results using DDAMFN as a feature extractor are practically the same as those obtained by fine-tuning the entire model. This similarity in performance can be attributed to the architecture of DDAMFN, which has been trained to generalize exceptionally well for the multitask challenge problem. This inherent capability allows it to extract meaningful features that are sufficient for achieving high performance without the need for extensive fine-tuning.
In contrast, the single-task experiment revealed a different scenario. Fine-tuning the entire model for the single-task yielded better results than merely training the classifier, and even outperformed the multitask results. This finding underscores the importance of task-specific optimization. When the model is fine-tuned for a specific task, it can leverage the nuances and particularities of that task, leading to superior performance.
This suggests that with meticulous optimization of the loss functions and careful consideration of data imbalance, the DDAMFN architecture could potentially achieve even better results in the multitask challenge. Proper handling of these aspects could unlock the full potential of the model, leading to significant performance gains in multitask settings.
Moreover, it is crucial to highlight the importance of effective threshold optimization in the Action Units task. By fine-tuning the thresholds, the results improved significantly, with an increase of 0.5 in performance metrics. This improvement demonstrates that beyond model architecture and training strategies, the post-processing steps such as threshold optimization play a vital role in achieving optimal results. Effective threshold optimization ensures that the model’s predictions are more accurate and reliable, contributing to overall performance enhancement.
In summary, while DDAMFN as a feature extractor performs on par with full model fine-tuning in a multitask setting, the single-task fine-tuning shows that there is room for improvement with targeted optimization strategies in multitask scenarios. This includes better loss function optimization, addressing data imbalance, and refining threshold settings, all of which are crucial for maximizing the performance of the model in the multitask challenge.